OpenStudy (anonymous):

solve : x^2 +y^2 = 41 x - y = 1

OpenStudy (anonymous):

y=5,4 and x= 5,6

OpenStudy (anonymous):

From second equation x = 1 + y Substituting in first equation (1 + y)^2 + y^2 = 41 1 + 2y + y^2 + y^2 = 41 2y^2 + 2y + 1 -41 = 0 2y^2 + 2y - 40 = 0 Solve this to get value of y substitute in second equation and get value of x

OpenStudy (anonymous):

x^2+y^2=41

OpenStudy (anonymous):

the answer is xy+2=41

OpenStudy (anonymous):

Can u pls read your question again and rewrite the full question

OpenStudy (shadowfiend):

Don't forget to give a medal if someone was helpful!